mathslover (mathslover):

IF: \(\large{\frac{x}{a}=\frac{y}{b}+\frac{z}{c}}\) Then show that : \[\large{\frac{x^3+a^3}{x^2+a^2}+\frac{y^3+b^3}{y^2+b^2}+\frac{z^3+c^3}{z^2+c^2}=\frac{(x+y+z)^3+(a+b+c)^3}{(x+y+z)^2+(a+b+c)^2}}\]

OpenStudy (anonymous):

Nice question.. Can't this be more difficult ??

mathslover (mathslover):

x/a = y/b = z/c = k I tried like this : @waterineyes :D

OpenStudy (anonymous):

There is plus sign in between or they all are equal ??

mathslover (mathslover):

there is plus

OpenStudy (anonymous):

Can you put then equal then ???

OpenStudy (anonymous):

*them

mathslover (mathslover):

nope

mathslover (mathslover):

How we will put equal @waterineyes ?: well .. this is what i did

OpenStudy (anonymous):

I am thinking what we can do in it..

mathslover (mathslover):

x=ak y=bk z=ck ?

mathslover (mathslover):

@vishweshshrimali5 @TuringTest @asnaseer @KingGeorge

OpenStudy (anonymous):

You cannot put equal they are in plus form..

mathslover (mathslover):

no no .. : x/a = y/b = z/c = k x=ak , y = bk and z = ck

OpenStudy (anonymous):

k = k + k No they are not equal..

mathslover (mathslover):

what? k = k + k ???

OpenStudy (anonymous):

You are letting them as k..

mathslover (mathslover):

yea : let x/a = y/b = z/c = k

OpenStudy (anonymous):

See I show you..

mathslover (mathslover):

k

OpenStudy (anonymous):

\[\large \frac{x}{a}=\frac{y}{b}+\frac{z}{c} \implies k \ne k + k \;\; ( Agree?)\]

mathslover (mathslover):

yes it implies that : k is not equal to 2k

OpenStudy (anonymous):

So how can you put them as k ??

OpenStudy (anonymous):

x/a is sum of y/b and z/c..

mathslover (mathslover):

Oh wait .. :from your side a slap to me :( mistake there: x/a = y/b = z/c is the correct situation of "if"

OpenStudy (anonymous):

If they were equal then you can put them as k..

mathslover (mathslover):

yes but wht next

OpenStudy (anonymous):

I am still thinking..

mathslover (mathslover):

ok

OpenStudy (anonymous):

We will need here more mathematicians.. @mukushla can you help here..

mathslover (mathslover):

right have to go now.. sorry and please think abt this more

OpenStudy (anonymous):

Yeah don't worry.. I am working on it..

OpenStudy (anonymous):

oh man thats equal not plus...i think we did it before...

OpenStudy (anonymous):

Really ??

OpenStudy (anonymous):

mathslover is offline.. We have to solve it somewhat..

OpenStudy (anonymous):

Yes they are equal..

OpenStudy (anonymous):

@mathslover you have solved half part Now solve for RHS. In RHS just replace x = ak, y = bk and z = ck After factoring you will get : \((a + b + c) \times \frac{k^3 + 1}{k^2 + 1}\) which is equal to LHS..

OpenStudy (anonymous):

yeah... and now we just need to put\[k=\frac{x+y+z}{a+b+c}\]

OpenStudy (anonymous):

But if we will solve Somewhat for LHS and somewhat for RHS then also we can prove them equal if they are coming equal too.. @mathslover your question should be actually like this: If \[\frac{x}{a} = \frac{y}{b} = \frac{c}{z}\] Then .......................

mathslover (mathslover):

that is what i am thinking then?

OpenStudy (anonymous):

I think you posted the same question..

mathslover (mathslover):

Yes but didn't get the answer

OpenStudy (anonymous):

How??

OpenStudy (anonymous):

You have done half part correctly..

OpenStudy (anonymous):

Now change the value of k according to mukushla gave above..

mathslover (mathslover):

OH k =(x+y+z)/(a+b+c) but i had taken k = x/a = y/b = z/c so how it will be k = (x+y+z)/(a+b+c)

OpenStudy (anonymous):

\[k = \frac{x}{a}=\frac{y}{b}=\frac{z}{c} \;\; then \;\; k = \frac{x + y + z}{a + b + c}\]

OpenStudy (anonymous):

For example: \[k= \frac{1}{2} = \frac{2}{4} = \frac{3}{6} \implies k = \frac{1 + 2 + 3}{2 + 4 + 6} \implies k = \frac{6}{12} = \frac{1}{2}\]

mathslover (mathslover):

right thanks a lot i got it .. thanks @mukushla and @waterineyes

OpenStudy (anonymous):

Welcome dear..

OpenStudy (anonymous):

let me complete the solution\[\frac{x}{a}=\frac{y}{b}=\frac{z}{c}=k\] so\[x=ak\]\[y=bk\]\[z=ck\]add them \[x+y+z=k(a+b+c)\]so\[k=\frac{x+y+z}{a+b+c}\]

OpenStudy (anonymous):

That makes sense in a better and understandable way..

mathslover (mathslover):

thanks @mukushla that helped

OpenStudy (anonymous):

np...man

OpenStudy (jiteshmeghwal9):

similar questions i also gt in my N.T.S.E. exam bt i gt confused whereas they were too easy to solve , HAHAHA :)

mathslover (mathslover):

that is the speciality of NTSe

OpenStudy (jiteshmeghwal9):

yes!

OpenStudy (anonymous):

\[ x=a k\\ y=b k\\ z= c k \\ \frac{a^3+x^3}{a^2+x^2}+\frac{b^3+y^3}{b^2+y^2}+\frac{c^3+z^3}{c^2+z^2}=\\ \frac{a^3 k^3+a^3}{a^2 k^2+a^2}+\frac{b^3 k^3+b^3}{b^2 k^2+b^2}+\frac{c^3 k^3+c^3}{c^2 k^2+c^2}=\\ \frac{\left(k^3+1\right) (a+b+c)}{k^2+1}\\ \] The other side \[ \frac{(a+b+c)^3+(x+y+z)^3}{(a+b+c)^2+(x+y+z)^2}=\\ \frac{(a k+b k+c k)^3+(a+b+c)^3}{(a k+b k+c k)^2+(a+b+c)^2}=\\ \frac{\left(k^3+1\right) (a+b+c)}{k^2+1} \]